provide the number in the box [ ], such that (2/3) × [ ] = (10/30)

Grade:12

3 Answers

Harshit Singh
3 years ago
Dear Student

Let the required number be x,
Then,
= (2/3) × (x) = (10/30)
By cross multiplication,
= x = (10/30) × (3/2)
= x = (10 × 3) / (30 × 2)
= x = (5 × 1) / (10 × 1)
= x = 5/10
∴The required number in the box is (5/20)

Thanks
Yatheesh Kumar.P
2 years ago
2/3=x=10/30
dear friends
x=10/30 divided by 2/3=5/10
so the x or [] is 5/10
this is a easy way
thank you
 
Yatheesh Kumar.P
2 years ago
(a) (i) provide the number in the box [ ], such that (2/3) × [ ] = (10/30)
Solution:- Let the required number be x,
Then, = (2/3) × (x) = (10/30)
By cross multiplication,
= x = (10/30) × (3/2)
= x = (10 × 3) / (30 × 2)
= x = (5 × 1) / (10 × 1)
= x = 5/10
∴The required number in the box is (5/20)